### v1

# MCGMark: An Encodable and Robust Online Watermark for LLM-Generated Malicious Code

Abstract
With the advent of large language models (LLMs), numerous software service providers (SSPs) are dedicated to developing LLMs customized for code generation tasks, such as CodeLlama and Copilot. However, these LLMs can be leveraged by attackers to create malicious software, which may pose potential threats to the software ecosystem. For example, they can automate the creation of advanced phishing malware. To address this issue, we first conduct an empirical study and design a prompt dataset, MCGTest, which involves approximately $400$ person-hours of work and consists of $406$ malicious code generation tasks. Utilizing this dataset, we propose MCGMark, the first robust, code structure-aware, and encodable watermarking approach to trace LLM-generated code. We embed encodable information by controlling the token selection and ensuring the output quality based on probabilistic outliers. Additionally, we enhance the robustness of the watermark by considering the structural features of malicious code, preventing the embedding of the watermark in easily modified positions, such as comments. We validate the effectiveness and robustness of MCGMark on the DeepSeek-Coder. MCGMark achieves an embedding success rate of $88.9\%$ within a maximum output limit of $400$ tokens. Furthermore, it also demonstrates strong robustness and has minimal impact on the quality of the output code. Our approach assists SSPs in tracing and holding responsible parties accountable for malicious code generated by LLMs.
Index Terms:
Traceability, Watermark, Large Language Models, Code Generation
1 Introduction
Code generation has become an important topic in software engineering [1], enabling the automatic generation of code snippets from the natural language requirements provided by users. This advancement significantly reduces the manual coding effort and simplifies the software development process [2]. Recently, with the advent and development of large language models (LLMs), their capacity for enhancing code-related tasks has been notably recognized [3]. Meanwhile, Software Service Providers (SSP) are dedicating efforts to develop LLMs specifically tailored for code generation tasks, such as CodeLlama [4] and DeepSeek-Coder [5].
Although LLMs offer remarkable capabilities for code generation, they are often exploited by attackers for misbehaving activities. Previous research [6] reveals that participants in online criminal forums are exploiting LLMs to develop malicious software, e.g., spyware and ransomware. Furthermore, reports from various organizations, such as Meta [7] and Cybersecurity [8], indicate a rising trend in malicious software and network attacks caused by LLMs. Numerous instances and posts also demonstrate the use of LLMs to generate malicious code on various technical forums [9, 10], which could potentially endanger users’ assets.
However, SSPs face significant challenges in tracing the sources of malicious code generated by LLMs [11]. Identifying the specific malicious developer who utilizes an LLM to generate a given piece of malicious code remains an unresolved challenge in the current landscape. This makes the malicious use of LLMs for code generation a low-risk but high-return activity. To mitigate this issue, SSPs like OpenAI have also proposed various solutions, such as zero-shot detectors [12] and fine-tuning language model detectors [13]. However, subsequent research has shown that these solutions are ineffective or have poor robustness [14]. For example, OpenAI’s classifier, which had an accuracy rate of only $26\%$, has now been discontinued by OpenAI [15].
Watermarking technology is a promising solution for tracing the origin of the content generated by LLMs [16, 17]. It involves incorporating explicit or hidden features, such as special lexical substitution, into the content to distinguish whether it is generated by LLMs or humans [18]. However, current watermarking methods may not meet SSP’s requirements for tracing the source of malicious code [19, 20]. Firstly, they suffer from poor traceability and implicitness. The most advanced LLM watermarking technologies can only identify whether a piece of text was generated by an LLM, but they cannot reflect the user’s identity [21, 22]. This limitation prevents SSP from tracing malicious developers. Additionally, most watermarking techniques are post-processing based [23, 24, 17], typically involving word substitution and statistical rules, which only modify the LLM’s generated output. Such watermarks often leave noticeable characteristics in the code, making them more easily detectable and modifiable by malicious developers [16]. Secondly, maintaining code generation quality is challenging. The most advanced implicit watermarking techniques primarily target text generation [25, 26], while code typically has a more rigid structure and stricter syntax rules [27]. Embedding implicit watermarks during code generation while preserving code quality presents a significant challenge. Lastly, robustness is a concern. Certain elements in the code, such as comments and variable names, can be modified without affecting the code’s functionality [28]. Malicious developers can easily alter these elements, making the watermark vulnerable to removal. Therefore, it is necessary to consider the structural characteristics of the code when embedding watermarks to ensure that they are not easily compromised.
To address the aforementioned issues, in this paper, we propose MCGMark, the first watermarking method specialized for tracing code generated by LLMs. Our approach implicitly embeds encodable watermarks during the code generation process, considering the code’s structure and ensuring the quality of the generated code. Firstly, MCGMark implicitly embeds encodable watermarks by controlling the LLM’s token selection, ensuring the watermark is difficult to discern while reflecting the generator’s identity. Secondly, MCGMark maintains the output quality of the LLM by handling the probabilistic outliers of candidate tokens. Lastly, we design a watermark skipping mechanism based on code structure and syntax rules, allowing MCGMark to decide whether to embed the watermark in subsequent code elements based on the code already generated. This ensures that watermarks are not embedded in easily modifiable code elements, such as comments and variable names, thereby enhancing the robustness of the watermark. Additionally, we conduct an empirical study on existing instances of malicious code. We collect $129$ real malicious code examples generated by LLMs and analyze $21,959$ malicious code repositories on GitHub. Based on the empirical study, we construct the first prompt dataset for LLM-generated malicious code, MCGTest, consisting of $406$ malicious code generation tasks to inspire watermark design and test its effectiveness.
MCGMark is designed as a plug-in, eliminating the need for additional models and data during both watermark embedding and detection processes. To implement MCGMark, the SSP only needs to adapt its token-matching rules to align with the specific LLM vocabulary in use. We adapt and apply MCGMark on DeepSeek-Coder-6.7b-instruct to evaluate its performance. MCGMark embed a total 24-bit watermark in $400$ tokens, achieving a watermark embedding success rate of $88.9\%$. Additionally, we conduct a user study involving $500$ program pairs and $1200$ modification attacks. The analysis demonstrates MCGMark’s effectiveness in maintaining code generation quality and resisting modification attacks.
In summary, this work contributes the following:
  • We construct the first malicious prompt dataset, MCGTest, consisting of $406$ prompts for LLM’s malicious code generation. based on MCGTest, we design and test the proposed watermarking method. This dataset also inspires further adversarial work against LLM’s malicious code generation.
  • We propose MCGMark, the first robust, code structure-aware, and encodable watermarking scheme to trace LLM-generated code. MCGMark implicitly embeds user identity information in code generation, ensuring both code quality and robustness against watermark tampering.
  • We apply MCGMark to DeepSeek-Coder-6.7b-instruct. The experimental results show that it successfully embeds a $24$-bit watermark and achieves a success rate of $88.9\%$ under a maximum LLM output of $400$ tokens. Additionally, MCGMark ensures code generation quality and withstands various modification attacks.
  • We publish the source code of MCGMark, and MCGTest at: https://github.com/KevinHeiwa/MCGTM
2 Background and Challenges
Figure 1: The motivation example and challenges of design watermark against LLM-generated malicious code.
2.1 Code Generation of LLM
Large Language Models (LLMs) are large-scale language models based on the Transformer architecture [29] and are trained on massive corpora, typically with billions of parameters or more [30]. In recent research, LLMs have demonstrated impressive performance in code generation tasks [31] , which can significantly improve software development efficiency [32]. The performance of LLMs in code generation tasks has received extensive research attention [33, 32]. Moreover, some Software Service Providers (SSPs) have trained/fine-tuned dedicated LLMs specifically for code generation, known as Code LLMs. Currently, SSPs have developed numerous popular Code LLMs, such as Code Llama [4], DeepSeek-Coder [5], and StarCoder2 [34].
2.2 Motivation
LLMs are increasingly being exploited by malicious developers for generating malicious code [35, 36, 37]. Numerous instances have demonstrated the high efficacy of LLMs in producing harmful software. Fig. 1 illustrates a real-world example where an LLM was prompted to generate code for stealing browser history  [^1] https://github.com/AI-Generated-Scripts/GPT-Malware. This irresponsible utilization of LLMs for malicious code generation could pose a significant threat to the security of the software ecosystem.
However, unlike malicious text, which often exhibits overt harmful features, certain code functionalities (e.g., keyboard logging) may not exhibit apparent malicious characteristics [38]. This subtlety renders LLMs vulnerable to exploitation for generating harmful code [39]. Furthermore, malicious developers can employ instruction hijacking [40] and jailbreaking [41] to further facilitate the generation of malicious code through LLMs. Therefore, it is imperative to design alternative approaches for LLMs to combat malicious code generation, with watermarking schemes emerging as one of the most promising solutions [20].
2.3 Challenges
Designing watermarks to trace the generation of malicious code introduces several novel challenges.
  • Traceability & Implicitness. The watermark must be implicit and accurately reflect the user’s ID. Fig.1.(1).a shows an example of a watermark from the literature [21], but applying this watermark only indicates whether the code was generated by an LLM, failing to trace back to a specific user. Fig.1. (1).b illustrates a pattern from the watermarking technique in the literature [23], which is a post-processing watermark. This technique does not participate in code generation, it processes the code after generation using rules such as code transformation [24]. However, this watermark depends on predefined code transformation patterns, limiting its adaptability and failing to ensure compatibility with all code variations. Furthermore, relying on predefined patterns makes the watermarked code exhibit noticeable characteristics, which can be easily detected and disrupted by malicious developers.
  • Ensuring Code Generation Quality. The embedding of watermarks must maintain the quality of the generated code. Fig.1.(2) shows an example of a hard watermark from the literature [25]. In this instance, watermark embedding significantly degraded the code generation quality, rendering the LLM-generated code unusable. Compared to text, code contexts typically exhibit lower entropy and follow more fixed patterns in generation [27]. Some multi-encoding watermarking techniques rely on the strong generative capabilities of LLMs to mitigate the impact of watermarking on generation quality [42, 43]. However, this approach is not suitable for code generation. Even minor perturbations in code can significantly affect its functionality and usability. Thus, during the LLM’s code generation process, minimizing the impact of watermark embedding on code quality is a crucial challenge.
  • Resistant to Tampering. Code elements, such as comments and variable names, can be altered without affecting the code’s functionality. If a watermark is embedded in these elements, it can be easily tampered with. Fig.1.(3).a demonstrates a watermarking scheme from the literature [19], where the watermark is added to variable names and can be easily modified. Fig.1. (3).b shows an example of a soft watermark from the literature [25], where the watermark is added to comments and can also be easily disrupted. However, during the online embedding of watermarks, the generation of code and the embedding of watermarks need to occur simultaneously. When embedding a watermark, we cannot access the complete code generated by the LLM. Thus, we need to make decisions about watermark embedding for the yet-to-be-generated code based on incomplete code, which presents a significant challenge.
3 MCGTest: A prompt dataset for LLM malicious code generation
In this section, we conduct an empirical study using real-world malicious code to help design our watermark. This study results in the creation of MCGTest, a comprehensive dataset of malicious code generation prompts that includes both actual instances of LLM-generated malicious code and potential scenarios.
3.1 Data Collection
To comprehensively collect malicious code generation scenarios with LLMs, we include both real examples and potential use cases.
Existing Instances Collection. This involves gathering existing instances of malicious code generated by LLMs. We collect data from two major technical communities, GitHub [44] and Stack Overflow [45], three main literature databases, Google Scholar [46], arXiv [47], DBLP [48], and the Google search engine [49]. We use the following four keywords for the above six platforms to collect results from January 2023 to March 2024: “large language model malicious code,” “large language model malware,” “GPT malware,” and “GPT malicious code.” In total, we collect $3,644$ results, including code repositories, papers, and articles.
Potential Scenarios Collection. This involves gathering possible scenarios for using LLMs to generate malicious code. For this part, we primarily collect repositories related to malicious code from GitHub. Using the keywords “malicious code” and “malware,” we identify and collect $21,959$ malicious code repositories.
3.2 Data Pre-processing
In this process, we describe the preprocessing of the collected data to extract malicious code instances, ensuring the relevance of the data for subsequent analysis.
To filter out real instances and potential scenarios of malicious code generation by LLMs, the data preprocessing process is divided into two steps, as shown in Fig. 2. First, we filter out extraneous data, such as empty code repositories and duplicate results. Second, we extract instances of malicious code from the filtered results to help us understand the functionality of malicious code and construct the MCGTest dataset.
To ensure the reliability of this analysis. Four researchers with more than four years of software development experience are assigned to filter and analyze the collected results. They are classified into two groups. Group 1 analyzes the data from Part 1, while Group 2 is responsible for the data of Part 2. Members in each group work independently and then align conflicting results. Each group focuses only on their respective portion of the data, without interfering with each other.
Figure 2: The process of data pre-processing.
3.2.1 Filtering Extraneous Data
In this step, we primarily filter out results unrelated to malicious code. Group 1 removes irrelevant data of Part 1, including empty code repositories, advertisements without substantial content, and web pages with only titles. Additionally, duplicate results and results unrelated to malicious code, such as evaluations or fixes using LLM for malicious code or security vulnerabilities of LLM itself, also be removed. Group 2 primarily selects high-quality malicious code repositories from Part 2, excluding repositories unrelated to the topic or invalid repositories, such as malicious code detection tools or repositories that do not provide access to source code directly. Furthermore, to ensure an adequate number of collected malicious code instances, repositories with at least $200$ stars will be considered [50]. After alignment within the groups, Group 1 obtains a total of $306$ results, including $128$ literature references, five code repositories, and $173$ relevant web pages and articles. Group 2 collects $93$ code repositories.
3.2.2 Extracting Malicious Code Instances
In this step, both groups are tasked with obtaining instances of malicious code generated by LLMs from the filtered results obtained in the previous stage. This forms the foundation of our LLM malicious code prompt dataset. Group 1 extracts descriptions of malicious code generated by LLMs from literature, news articles, and code repositories. They also identify and analyze malicious code snippets to determine their functionality. Group 2, on the other hand, selects usable malicious code functionality from the malicious code repositories. They exclude incomplete or ambiguous code and functions that are unrelated to the repository description or purpose, such as data visualization or graphical user interfaces. Additionally, to maintain the quality of collected instances, Group 2 members will exclude functions with fewer than five lines of code. After alignment within the groups, Group 1 and Group 2 have selected $129$ and $395$ instances of malicious code, respectively.
In summary, we collect a total of $524$ malicious code instances, encompassing both LLM-generated samples and a comprehensive survey of prominent open-source malicious code. thereby encompassing potential scenarios for malicious code generation.
3.3 The Construction Process of Prompt Dataset
In this process, we construct a prompt dataset called MCGTest for LLM malicious code generation, based on the data collected from Section 3.2.
Figure 3: The construction process of MCGTest.
Fig. 3 illustrates the construction process of MCGTest. This process contains three steps. First, we create functional summaries of the malicious code instances extracted in Section 3.2. Next, we filter these instances to exclude those that do not meet our requirements, such as redundant functionalities. Finally, we use the filtered set of malicious code functionalities to construct a dataset of prompts for malicious code generation.
3.3.1 Summary of Malicious Code Functionality
In this step, we aim to collect comprehensive information about malicious code’s functionality, including the overall malicious purpose of each instance and the specific malicious functions within them. To achieve this, we establish a Malicious Function Set (MFS) to compile all these functionalities. First, we summarize the functionality of each collected malicious code instance based on descriptions from their respective repositories or literature and add these summaries to the MFS. Next, we divide all instances into individual functions and use GPT-4 [51], an advanced LLM, to generate code summaries for these functions. Finally, we incorporate these summaries into the MFS. At this point, the MFS contains both the overall malicious purposes of all instances and the specific malicious functions found within them, ensuring a thorough representation of malicious functionalities at both the instance and function levels within our dataset.
3.3.2 Filtering Malicious Functionalities
In this step, we filter malicious functionalities, since not all of them are malicious in MFS. For example, some serve as copyright declarations. Additionally, there may be redundancy in functionalities across different functions. Therefore, we need to filter out malicious code functionalities from MFS. During the filtering process, we employ a closed card sorting method to ensure the accuracy of results. Closed card sorting is one of the most efficient methods for organizing information into logical groups [52]. Two participants are involved; both have over four years of programming experience. The card title is code functionalities, and the description consists of code or descriptions from the code repository/literature. Both participants read and filter cards according to specified criteria, aligning their results. The overall filtering rules are as follows:
  • Remove cards with duplicate semantic titles;
  • Remove cards with ambiguous semantic titles, such as functions with unclear meanings;
  • Remove cards with non-malicious semantic titles, such as copyright declaration functions.
Following these rules, we obtain a total of $406$ cards, comprising $72$ from Part 1 and $369$ from Part 2. These $406$ cards correspond to $406$ distinct malicious code functionalities.
3.3.3 Creating the Malicious Prompt Dataset
In this step, we construct the MCGTest dataset based on $406$ filtered cards. We primarily use the titles on each card describing the malicious code functionality to craft corresponding prompts. As a result, MCGTest comprises $406$ prompts. The prompt creation task involved three participants, each with an average of over four years of programming experience. Initially, each prompt was assigned to two participants: one for writing and one for review. In cases of disagreement, the third participant would facilitate a discussion to reach a consensus. For the prompt format, we drew inspiration from MBPP and HumanEval [53], two well-known datasets for evaluating LLM code generation performance. Each prompt consists of a function description and a function name, as illustrated in Fig. 4. MCGTest is designed to be compatible with various LLMs. Typically, an LLM has multiple versions, differing mainly in parameter count and whether they are Base or Chat versions [27, 54]. Base versions are usually continuation models with limited human interaction capabilities, while Chat versions can engage in dialogue. MCGTest’s prompt format is compatible with both versions.
In summary, we construct $406$ malicious code generation tasks for MCGTest. These tasks include real instances of LLM-generated malicious code as well as prominent potential scenarios. Additionally, MCGTest is compatible with LLMs and can be easily accessed from the open-source repository referenced in this paper. Beyond designing our watermarking scheme based on MCGTest, it will inspire further research on LLM-generated malicious code.
Figure 4: The prompt format of MCGTest.
4 MCGMark:An Encodable and Robust online watermark for LLM code generation
In this section, we introduce MCGMark, a method for embedding encodable watermarks over LLM code generation. MCGMark allows us to trace the developer of the generated malicious code. The watermark is implicit, and its robustness is enhanced by leveraging the structural features of the malicious code, making it difficult to be aware and resistant to tampering.
Figure 5: The overview of watermark embedding and detection.
4.1 Overview
Fig. 5 outlines the watermark embedding and detection process of MCGMark. Each process comprises five steps. It is noteworthy that both the embedding and detection of the watermark do not require the introduction of additional models or external data. Furthermore, during the watermark detection phase, there is no need to reload the LLM. This aligns well with the practical needs of SSPs.
For the watermark embedding process. MCGMark first initializes the watermark based on the user’s ID. The watermark consists of detection bits and error-correction bits, which collectively represent the user’s ID (Section 4.3.1 and Section 4.5.1). Subsequently, MCGMark partitions the LLM’s vocabulary and embeds multi-encoded watermarks by controlling the LLM’s token selection (Section 4.2). To mitigate the impact on code generation quality during watermark embedding, MCGMark then processes probability outliers in the vocabulary to ensure the LLM selects high-probability tokens, and updates the error-correction bit information (Section 4.3). Next, to enhance watermark robustness, we design a watermark skipping strategy for MCGMark based on code structure and syntax. MCGMark implements this strategy based on the generated code elements, ensuring that watermarks are not added to easily modifiable code elements. Finally, as code generation progresses, watermarks are embedded in a round-robin fashion to further enhance their robustness (Section 4.4).
For the watermark detection process. MCGMark first tokenizes the code into a sequence of tokens with the tokenizer. Next, MCGMark removes tokens from the sequence that would have been skipped during the embedding process, based on the employed skip strategy. Subsequently, MCGMark partitions the vocabulary and examines the vocabulary membership of tokens in the sequence to recover the watermark information. Since the watermark is embedded in multiple rounds, MCGMark then trims the recovered watermark. Finally, MCGMark reconstructs the user’s ID from the multiple segments of the trimmed watermark (Section 4.5.2).
4.2 Encodable Watermark Embedding
In this process, MCGMark embeds encodable watermarks during the code generation by controlling token selection.
4.2.1 LLM Code Generation Process
LLM Code Generation Process serves as the foundation for watermark embedding in MCGMark. In a typical LLM code generation process, the LLM maintains a token-level vocabulary $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$, typically comprising approximately $3.2\times 10^{4}$ tokens (e.g., the DeepSeek-Coder-6.7b utilizes a vocabulary of $32,022$ tokens) [55]. When a prompt $R$ is input into the model $M$, it first employs a tokenizer to segment $R$ into token-level components, $R\Rightarrow T=\left\{t_{0},t_{1},\cdots,t_{m}\right\}$. Subsequently, $M$ computes the probability distribution $P_{1}=\left\{P_{0}^{\prime},P_{1}^{\prime},\cdots,P_{n}^{\prime}\right\}$ for all tokens in the vocabulary based on $T$. Then, $M$ selects the highest probability token $v_{i}^{\prime}(0\leqslant i\leqslant n)$ from the vocabulary as the generated token and appends it to the prompt $R$. This process transforms the prompt $R$ into $R^{\prime}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime}\right\}$. $R^{\prime}$ is then fed back into model $M$, and this process iterates until a predetermined generation length $L$ is reached. The final output of the model is represented as $R^{(L)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(L)}\right\}$, where $\left\{v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(L)}\right\}$ constitutes the generated code [56].
4.2.2 Watermark Embedding
This step delineates the watermark embedding process in MCGMark. Inspired by the study [25], MCGMark incorporates watermarking by modifying the vocabulary $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$. MCGMark first generates a random set $D$ ($0<|D|\leqslant|V|$) with a hash value $H$. The elements in $D$ are unique, increasing integer values representing positions in the vocabulary. These elements form the selected vocabulary $\mathbb{A}$, while the remaining elements constitute vocabulary $\mathbb{B}$. MCGMark controls the LLM’s token selection by modifying the probability distribution of the vocabulary, constraining it to choose from the selected vocabulary. Simultaneously, to prevent excessive dependence on $H$ in vocabulary partitioning, MCGMark applies a pseudo-random augmentation to $H$. This approach ensures that $H$ remains variable while maintaining it reproducible. As all generated tokens stem from randomly chosen vocabulary subsets, the LLM-generated code inherently differs from manually written code. The probability of complete overlap between manually-generated and model-generated code is merely $\frac{1}{2^{L}}$ [25]. This low probability ensures the effectiveness of the code watermark.
4.2.3 Encoding Watermark
MCGMark encodes the watermark information to represent user-specific data. Specifically, MCGMark achieves encoding watermark embedding by modifying the probabilities of the LLM vocabulary. For embedding watermark $w_{w}$ into token $v_{i}^{(v)}$, MCGMark controls the LLM’s selection as follows:
  • If the watermark bit is $1$, select a token from $\mathbb{A}$.
  • If the watermark bit is 00, select a token from $\mathbb{B}$.
To ensure watermark correctness, MCGMark must guarantee that the model selects elements from the predefined vocabulary. This requires that the token with the highest probability in the modified vocabulary is present in the selected vocabulary. After generating element $v_{i}^{(v-1)}$, MCGMark obtains the probabilities of all tokens in the current vocabulary and calculates: $P_{\text{gap }}=\max\left\{P_{v}\right\}-\min\left\{P_{v}\right\}~{}.$
MCGMark then adds $P_{\text{gap}}$ to all elements in the selected vocabulary to ensure this vocabulary contains the highest probability token.
Thus, MCGMark completes the online multi-encoding watermark embedding in code generation.
4.3 Ensuring Watermarked Code Quality
In this process, MCG mitigates the impact of watermarking on code quality by addressing probability outliers in the vocabulary.
4.3.1 Code Quality Enhancement with Outliers
To minimize the impact of watermarks on LLM code generation quality and process high-probability tokens of the vocabulary, MCGMark develops a watermark quality enhancement algorithm based on the probability outlier values of the LLM’s vocabulary, as shown in Algorithm 1. This algorithm addresses potential issues arising from vocabulary partitioning, which may lead to incorrect selection of deterministic tokens and propagate errors in subsequent code generation. Leveraging the powerful generation capability of LLM, MCGMark implements the following strategy to ensure code generation quality during watermark embedding.
  1. 1.
    Before generating token $v_{i}^{(v)}$, MCGMark analyzes the probabilities $P$ to identify upper outliers—tokens with significantly higher probabilities.
  2. 2.
    If no upper outliers exist, MCGMark proceeds with standard watermark embedding.
  3. 3.
    If upper outliers are present and only for a single outlier, MCGMark includes it into the selected vocabulary.
  4. 4.
    For multiple outliers, MCGMark randomly selects half with $H$ and includes them in the selected vocabulary.
However, outliers can affect the accuracy of watermark detection. During detection, we can only analyze the code and cannot obtain the real-time probability distribution of the vocabulary. This limitation aligns with real-world scenarios. To address this, we include error-correction bits in the watermark to recover the watermark information. When outliers impact the watermark information, MCGMark sets the error-correction bit to $1$; otherwise, it is set to 00. Once the watermark detection bits are fully embedded, the error-correction bits are also generated. Subsequently, the error-correction bits are embedded into the code. During the embedding of error-correction bits, the watermark is not influenced by outliers. This ensures accurate watermark detection while reducing the impact of the watermark on the embedded elements by more than $50\%$.
4.3.2 Outliers Detection
MCGMark employs the Inter-Quartile Range (IQR) method for outlier detection, based on the boxplot approach [57]. The IQR method is a robust statistical technique for describing data dispersion, measuring the distribution range of the middle $50\%$ of data by calculating the difference between the third and first quartiles of a dataset [58]. This method is particularly suitable for MCGMark due to its robustness to extreme values and its effectiveness in identifying outliers in non-normally distributed data, which is common in token probability distributions [59, 60]. MCGMark defines the upper whisker of the boxplot as:
$\small F_{\text{upper}}=Q_{3}+S\cdot IQR=(S+1)\cdot Q_{3}-S\cdot Q_{1}$ (1)
where $Q_{1}$ is the lower quartile, $Q_{3}$ is the upper quartile, $IQR=Q_{3}-Q_{1}$, and $S$ is a scaling factor. This approach allows us to identify and appropriately handle tokens that are crucial for maintaining code generation quality.
In summary, after this step, MCGMark achieves multi-encoding watermark embedding during LLM code generation while maintaining code quality.
Algorithm 1 Preserving LLM Code Generation Quality through Outlier Management
Input: Prompt: $R$, watermark: $W$, the number of new tokens: $L$, threshold for $P_{\text{dis }}$: $Thr\_P_{\text{dis }}$, a hash key $H$.
1:  for $l=0,1,\cdots,L$ do
2:     Obtain $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$ and $P_{l}=\left\{P_{0}^{l},P_{1}^{l},\cdots,P_{n}^{l}\right\}$.
3:     Calculate the outliers of $P_{l}$ with Equ. (1).
4:     if $F_{\text{upper }}\neq\emptyset$ then
5:        Case 1: $|F_{\text{upper }}|=1$,$D^{{}^{\prime}}=D\cup F_{\text{upper }}$.
6:        Case 2: $|F_{\text{upper }}|\geq 2$, $D^{{}^{\prime}}=D\cup F_{\text{upper }}[0:\lceil\frac{F_{\text{upper }}}{2}\rceil]$ with $H$.
7:     end if
8:     Sample $v_{i}^{(l)}$ from $V$.
9:     if $v_{i}^{(l)}\in F_{\text{upper}}~{}\&~{}v_{i}^{(l)}\notin D$ then
10:        Setting the watermark’s error-correction bit to $1$.
11:     end if
12:  end for
4.4 Enhancing Watermark Robustness
In this process, we provide a detailed description of MCGMark’s robustness scheme, designed based on the structure of malicious code. This scheme enables watermarks to withstand simple code modifications attempted by malicious developers. We delineate the adversarial scenarios and present a comprehensive overview of the design process.
4.4.1 Adversarial Scenarios
Malicious developers utilize LLM to generate malicious code. When these developers become aware of the potential presence of a watermark in the code, they may modify it. To avoid compromising the code’s execution, malicious developers tend to modify easily changeable elements, especially for less experienced “script kiddies.” Consequently, if the watermark is placed on easily modifiable elements, such as variable names, the watermark becomes highly susceptible to being rendered ineffective.
4.4.2 The Overall of Watermark Robustness Enhancement
This step presents the robustness enhancement scheme designed by MCGMark to address adversarial scenarios. Enhancing the robustness of watermarking in LLM-generated code involves skipping embedding watermarks in parts of the code that are easily modified or removed. However, since code generation and watermark embedding need to be synchronized, MCGMark must decide online whether to skip watermark embedding for the next token. To achieve this, MCGMark divides the process into two steps.
  1. 1.
    MCGMark determines which elements in the code should be skipped for watermark embedding. This is done by defining watermark skipping rules based on code structure, code cloning, and observations from malicious code instances.
  2. 2.
    MCGMark decides whether to embed a watermark for the next token, $v_{i}^{(l+1)}(l\in[0,L-1])$, based on the tokens already generated by the model, $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$. Since the tokens in the LLM vocabulary do not strictly adhere to human grammar rules, especially in the case of Code LLM, such as “:(” or “])”, we design a watermark skipping scheme for MCGMark based on the grammar rules of the code and the LLM vocabulary.
It is worth noting that due to significant differences in code structure among different programming languages, and since Python is currently one of the most commonly used languages by malicious developers [61], MCGMark focuses solely on Python language.
4.4.3 Watermark Skipping Rules
In this step, we describe the element selection criteria of MCGMark for watermark skipping. To design watermark skipping rules, it is crucial to identify which elements in Python code are easily modifiable without affecting code usability. Our analysis focuses on three key perspectives:
  • (Code Structure.) Python primarily consists of the following elements [62]: indentation, keywords, identifiers, statements, expressions, functions, objects, object methods, types, numbers, operators, comments, exception handling, input/output, and blank lines. Previous research indicates that modifications to certain elements have minimal impact on code execution quality: identifiers (including variable names, function names, and class names), comments, output statements, numerical values, and blank lines [63, 64].
  • (Code Clone.) Code clone focusing on identifying instances of code plagiarism using similarity metrics [65, 66]. And it’s classification typically into four levels: exact, lexical, syntactical, and semantic. Among these, exact clones are identical except for whitespace, layout, and comments. Lexical clones differ in identifiers but maintain the exact same structure. Syntactical clones involve the addition, modification, or deletion of statements while preserving the overall structure. Semantic clones, on the other hand, refer to code segments that achieve the same functionality but use different syntax. Recent studies indicate that detecting complete and lexical similarity is feasible [67, 68, 69, 70, 71, 72, 73], suggesting that modifying elements like whitespace, layout, comments, and identifiers is more manageable [74, 75, 76].
  • (Malicious Code Instances.) We analyze the existing instances of malicious LLM code in Section 3.2. We observe that assignments, comparisons, and parenthetical elements, besides comments, output, and identifiers, are also readily modifiable in these instances. Thus, when designing the watermark, we must avoid embedding it in elements related to these operations.
In summary, the watermark is not embedded in code elements listed in Table LABEL:tab:EazyModi.
4.4.4 Watermark Skipping Pattern
In this step, we design watermark skip patterns for MCGMark. The watermark embedding skip algorithm in MCGMark is based on watermark skipping rules. Since LLM-generated tokens are irreversible and the watermark embedding process is synchronized with the token generation, MCGMark cannot wait for the model to generate elements in Table LABEL:tab:EazyModi before skipping the watermark. The embedding modifies the distribution of $P_{l}=\left\{P_{0}^{l},P_{1}^{l},\cdots,P_{n}^{l}\right\}(l\in[0,L])$, influencing the LLM’s decision-making. Therefore, watermark embedding must be controlled before generating elements in Table LABEL:tab:EazyModi. Thus, MCGMark decides on watermark embedding for the next token based on the already generated code, considering that vocabulary tokens are irregular and may not correspond directly to code elements.
Seven patterns are designed for skipping watermark embedding during LLM code generation.
[…]
  • (Pattern 2.) If $v_{i}^{(l)}$ belongs to the set ${}\widehat{B}=\left\{(,[,^{\prime},",\{\right\}$, then no watermark is applied to subsequent tokens until a matching symbol is encountered. This is because elements in set $\widehat{B}$ are often followed by tokens containing identifiers, values, and other easily modifiable elements. Hence, MCGMark does not apply a watermark to the content inside parentheses or quotation marks. No processing is performed if a pair of matching symbols, such as “((((”,“))))”, appears within a single token.
[…]
  • (Pattern 4.) If $v_{i}^{(l)}$ is in the set $\widehat{D}=\left\{``````,^{\prime\prime},```\right\}$, representing multi-line comments, no watermark is applied to subsequent tokens until the same element reappears. Additionally, the watermark is rolled back by $1$ position.
[…]
  • (Pattern 6.) When one Pattern is in effect, another Pattern may be triggered, leading to conflicts in Pattern decisions. In such cases, if any Pattern is already active, subsequent Patterns will not be triggered. However, if conditions in $v_{i}^{(l)}$ satisfy the triggering criteria of two Patterns simultaneously, only the first Pattern in sequence will be triggered.
  • (Pattern 7.) Once all watermark bits have been added, but not all tokens have been generated, MCGMark continue iterating the embedding process to further strengthen the watermark.
4.4.5 Watermark Skipping Process
Algorithm 2 Enhancing the Robustness of LLM Online Code Watermark via Code Structure and Syntax
Input: Existing tokens: $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$ , watermark to be embedded $w_{X}$, set $\widehat{A,B,C,D}$.
1:  if $l\leq L$ then
[…]
3:     if not $\mathcal{LOCK}$ then
4:        if $v_{i}^{(l)}\in\left\{\widehat{A}\cup\widehat{B}\cup\widehat{C}\cup\widehat{D}\right\}$ then
[…]
7:           $V=V$, break.
8:        else
9:           if $X\leq x$ then
10:              Take set $D$ or $V\cap D$ corresponding to $w_{X}$, to $V$, $X=X+1$.
11:           else
[…]
13:           end if
14:        end if
15:     end if
16:     Based on the effectiveness of Pattern to determine $\mathcal{LOCK}\Leftarrow 0$.
17:     $V=V$, break.
18:  end if
In this step, we describe MCGMark’s watermark skip decision process based on the established skip rules and patterns. Algorithm 2 delineates the execution process of watermark skipping. After obtaining the sequence $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$, MCGMark first verify if $l\leq L$, where $L$ represents the maximum output token limit of the LLM. If this condition is met, MCGMark proceeds with the watermarking process; otherwise, MCGMark terminates. MCGMark then evaluates $v_{i}^{(l)}$ against several conditions: if it consists solely of whitespace characters, MCGMark triggers Pattern 5; MCGMark checks for any active patterns based on Pattern 6 which would preclude watermarking the next token; and MCGMark determines if $v_{i}^{(l)}\in\left\{\widehat{A}\cup\widehat{B}\cup\widehat{C}\cup\widehat{D}\right\}$, which would trigger the corresponding Pattern (1, 2, 3, or 4) along with Pattern 6 to prevent multi-pattern conflicts. If no pattern is triggered, MCGMark selects a word from vocabulary $\mathbb{A}$ or $\mathbb{B}$ based on the watermark information. Upon complete embedding of all watermark information, MCGMark triggers Pattern 7 to iterate and incorporate additional watermark data. Algorithm 2 ensures judicious watermark embedding while preserving code integrity and functionality, accounting for various code elements and potential pattern conflicts.
In summary, after these steps, the watermarks embedded by MCGMark are imperceptible and encode the code generator’s personal information. This approach simultaneously ensures code quality and enhances robustness by leveraging code structure.
4.5 Design and Detection of Watermark
In this process, we design watermark patterns for MCGMark to enhance watermark detection success rates and describe a lightweight watermark detection procedure.
4.5.1 Watermark Design
This step delineates the watermark prompt design for MCGMark. MCGMark employs a dual-component watermark design comprising Detection Bits and Error Correction Bits. Detection Bits primarily encode user information for traceability, while Error Correction Bits facilitate the recovery of potentially erroneous information in the Detection Bits. Although Algorithm 1 ensures LLM code generation quality, setting a low outlier threshold (small $S$ value) may result in excessive outliers, potentially impacting the LLM’s word selection. For instance, the LLM might be compelled to select words from vocabulary $\mathbb{B}$ instead of $\mathbb{A}$ as dictated by the watermark information bit, due to outlier presence. This scenario could lead to errors in specific watermark bits. To mitigate this issue, MCGMark introduces Error Correction Bits to restore information in the Detection Bits and generates watermarks in the Detection Bits without outlier influence. Furthermore, Detection Bits and Error Correction Bits are designed with equal length. This design effectively addresses the trade-off between maintaining code quality and preserving watermark integrity, enhancing the overall robustness of MCGMark’s watermarking strategy.
4.5.2 Watermark Detection
This step describes the lightweight watermark detection process for MCGMark. In detecting watermarks within code, MCGMark requires the code to be inspected, the model’s vocabulary, tokenizer, Algorithm 1, and Algorithm 2. There is no need to load an additional LLM. Given the malicious code, MCGMark first tokenizes the code using the tokenizer. Next, it applies the seven patterns and Algorithm 2 to remove elements where watermark embedding can be skipped, resulting in a sequence of code elements. Then, the hash value $H$ in Algorithm 1 is used to partition the vocabulary. The code element sequence is then traversed, and elements are categorized into the corresponding vocabulary parts, producing a sequence of 00s and $1$s. MCGMark subsequently segments this sequence according to the watermark length. Each segment provides detection bits and error-correction bits, which are used to obtain the user’s identity information using the following formula:
$\text{Detection Bits}\oplus(\text{1}~{}\&~{}\text{Error Correction Bits})~{}.$ (2)
Since our watermark is embedded in a round-robin manner, multiple segments of the watermark can sometimes be extracted from the code. When multiple rounds of watermarking are possible, consistent results from at least two rounds can help preliminarily identify the malicious code generator, thus enhancing fault tolerance.
5 Experiments
We evaluate MCGMark in this section based on the watermark design requirements analyzed in Section 2.3 with MCGTest. Specifically, we focus on the following research questions:
  • RQ1. What are the watermark embedding and detection success rates of MCGMark?
  • RQ2. How does MCGMark affect the quality of generated code?
  • RQ3. How robust is MCGMark in withstanding adversarial scenarios?
  • RQ4. What factors influence the successful watermark embedding in MCGMark?
For RQ1, we evaluate the watermark embedding success rate and detection success rate of MCGMark across 406 prompts included in MCGTest. For RQ2, we conduct a user study to assess the impact of watermark embedding on the quality of generated code. For RQ3, we design watermark attack patterns based on adversarial scenarios and evaluate the robustness of the watermarks. For RQ4, we primarily examine the influence of key parameters on MCGMark’s embedding success rate.
5.1 Experiment Setup
Evaluation Dataset: We use the MCGTest to validate the watermarking scheme. This dataset consists of $406$ malicious code prompts, including real instances generated by LLM and malicious code from high-quality open-source repositories.
LLM Selection: We test our watermarking strategy on the widely known Code LLM, DeepSeek-Coder-6.7b-instruct. This is because this LLM is open-source, allowing us to access its vocabulary (including $32,022$ tokens) and adapt MCGMark accordingly. Additionally, it has demonstrated high performance with relatively small parameters in multiple benchmarks [5].
Parameter: We set the number of LLM maximum output token, $L$, to $400$. Additionally, we set $\lceil\frac{|D|}{|V|}\rceil=0.5$. The total length of the watermark, $X$, is set to $24$ bits. And the watermark information is randomly generated. $Q3$ is set to $0.75$, and $Q1$ is set to $0.25$, which are commonly used settings for calculating outliers with quartiles [57].
Environment: All experiments are conduct on a workstation with $128$ CPU cores and $8$$\times$ NVIDIA A800 (80G) GPUs.
Comparison Tools: We use WLLM, the state-of-the-art LLM watermarking technique [25], as our baseline for comparison. WLLM is an open-source LLM online watermarking techniques, aligning closely with the objectives of MCGMark. Most other post-processing watermarking techniques are less suitable for code tasks and tend to impose fixed patterns on the code, making it more detectable and prone to tampering [14]. Therefore, we choose WLLM as a comparison benchmark. We adhere to the default parameter settings as specified in the paper, with $\gamma$ set to $0.25$ and $\delta$ set to $2$. To ensure fairness, the number of tokens generated by WLLM is also limited to $400$, and the same hash value as used in MCGMark is applied.
5.2 RQ1: Effectiveness of MCGMark
To answer RQ1, we evaluate the embedding success rate and detection success rate of MCGMark under a smaller token limit using the MCGTest dataset.
5.2.1 Watermark Embedding Success Rate
In the MCGTest dataset, MCGMark successfully embed watermarks in $361$ out of $406$ prompts, achieving an embedding success rate of approximately $88.9\%$. In contrast, WLLM successfully embed watermarks in $342$ prompts, with an embedding success rate of around $84.2\%$. MCGMark’s embedding success rate is slightly higher than WLLM’s, MCGMark’s watermarks can trace user information, while WLLM’s can only identify whether the watermark is generated by an LLM, failing to trace the source of malicious code.
We subsequently analyze the $45$ instances where MCGMark failed to embed watermarks. We find that $17$ tasks failed due to the generated code being too short to embed the watermark successfully. Additionally, $25$ tasks failed because the generated code triggered numerous traits, resulting in a high number of assignment statements and comments, which hindered watermark embedding. Another three prompts reject the malicious code generation request, resulting in empty content. Higher folding watermark encoding could potentially solve the issue of shortcode generation failures, though this is not the focus of this paper. The failures related to code structure primarily stemmed from our restriction of LLMs to generate a maximum of $400$ tokens. In practical applications, LLMs typically generate 20484096204840962048–40962048 – 4096 tokens or even more (e.g., DeepSeek-Coder-6.7b can generate up to 64K [5]). We retest the $25$ tasks that fail due to code structure with a maximum length setting of $2048$ tokens and successfully embed watermarks in $21$ of them, achieving an embedding success rate of $84\%$. Therefore, under conditions where token numbers are unrestricted or have a sufficiently high upper limit, MCGMark’s watermark embedding success rate could reach approximately $94.1\%$ across $406$ tasks.
5.2.2 False Rate Analysis
Out of the $361$ tasks where MCGMark successfully embed watermarks, $353$ watermarks are detected successfully, resulting in a detection success rate of approximately $97.8\%$. In contrast, WLLM does not support false positive rate checking. The eight instances where MCGMark failed to detect watermarks can be attributed to discrepancies in token splitting by the tokenizer. This issue leads to errors when verifying tokens against the vocabulary, returning incorrect results. This limitation is inherent to SSP, and we cannot improve the detection success rate by modifying MCGMark.
Answer to RQ 1: Under the $400$ token limit, MCGMark achieves a watermark embedding success rate of approximately $88.9\%$, with a detection success rate of around $97.8\%$. This demonstrates the effectiveness of MCGMark.
5.3 RQ2: Impact on LLM Code Generation
To answer RQ 2, we conduct a user study to evaluate the impact of MCGMark on the quality of LLM-generated. We primarily assess the code generation quality of the LLM to explore the impact of MCGMark on the code generation quality. In this part, we randomly select $50$ tasks from the $346$ successfully embedded tasks of MCGTest. Furthermore, we obtain $100$ code segments, with $50$ generated by the model using our watermarking strategy and $50$ without. We invite $10$ developers with at least $4$ years of development experience (excluding co-authors), including $6$ Ph.D. students, $2$ undergraduate students, and $2$ software engineers specializing in computer-related fields, to participate in our evaluation. We randomly shuffle the order of the $50$ code pairs and further shuffle the code order within each pair. We ask the developers to identify the code they believe contains a watermark for each code pair. We collect a total of $1000$ valid responses and organize the accuracy of these $1000$ responses, as shown in Table LABEL:table:rq2dis.
confidence interval $\approx\frac{\sum\text{Results}}{\text{NO. of participants}}\pm(t\times\frac{\text{Standard deviation}}{\sqrt{\text{NO. of participants}}})$ (3)
The recognition accuracy of $10$ participants for $500$ pairs of watermark/unwatermark code, totaling $1000$ segments, is $47.8\%$, which is close to random sampling. Moreover, the independent recognition rates of the $10$ participants, excluding the highest value (participant 5) and the lowest value (participant 6), fluctuate between $40\%$ to $50\%$. Furthermore, all these rates fall within the $95\%$ confidence interval (with Equ.(3), which is $[21.972,27.828]$). So, we can conclude that experienced practitioners with long-term development experience cannot correctly distinguish between watermark and unwatermark codes. This also demonstrates the stealthiness of our watermark and further confirms that the impact of MCGMark on code quality can be considered negligible.
Answer to RQ 2: MCGMark ensures the quality of the LLM’s code generation while embedding watermarks, without significantly affecting the model’s normal use.
TABLE III: Types and descriptions of attacks.
Types Attacks Description
Type 1 (1) modify/remove identifiers Modifying or removing variable names, function names, and class names, etc.
(2) modify/remove inputs and outputs Modifying or removing output and input content.
(3) modify/remove comments Modifying or removing single-/multi-line comments.
(4) modify/remove User-defined data Modifying or removing user-customizable data, such as numbers, strings, etc.
(5) modify/remove assignment statements Modifying or removing assignment operations, such as conditional statements and comparisons.
Type 2 (6) add comments Adding comments at any position.
(7) add assignment operation Adding assignment statements
(8) add redundant statements Adding useless statements, example of defining variables
5.4 RQ3: Resistance to Tampering
In this RQ, our main focus is on the robustness of the watermark in Section 4.4. Based on the literature [77], we primarily consider the following two types of attacks, as shown in Table III. These $8$ attack types cover the majority of modifications that typically employ against code watermarks [75, 70]. We also evaluate the performance of $50$ successfully watermarked codes under $8$ attacks, particularly whether these attacks affect the embedded watermark elements. For each attack, we conduct three attack instances. We carry out a total of $1200$ attacks.
In $1200$ attacks, we achieve a complete defense against Attacks 3, 6, 7, and 8. However, Attacks 4 and 1 failed $13$ times and $9$ times, respectively, resulting in defense success rates of $91.3\%$ and $94\%$. Upon carefully examining the failed instances, we identify a flaw in our token matching strategy during the implementation of Pattern 1-7. We rely on a generic regular expression matching approach, which proves inadequate for matching the tokens generated by LLMs due to their deviation from human language rules. The same phenomenon also affects the defense against Attack 2, resulting in an accuracy rate of $93.3\%$. Fortunately, this issue can be resolved by designing a more powerful matching scheme or SSP adapting the token vocabulary for the watermark.
The effectiveness of defense is comparatively reduced against Attack 5. Out of $150$ attacks of this type, there are $31$ defense failures, resulting in a success rate of $79.3\%$. We carefully examine these instances and find that $6$ of the failures were also due to matching issues. The remaining failures occur because of the maximum output limit of $400$ tokens. In such cases, the watermarking process could not be fully completed before reaching the maximum output limit. Fortunately, our watermarking design involves multiple rounds of watermark embedding, aiming to add as many watermarks as possible. As long as the watermark is added more than once, such cases have minimal impact on our detection performance. Additionally, relaxing the constraint on the maximum number of output tokens can also effectively address this issue.
Additionally, it is worth emphasizing that malicious developers may attempt to modify the code generated by other LLMs. However, as long as the generated LLM incorporates our watermark, the watermark information cannot be removed from the code.
Answer to RQ 3: MCGMark maintains a defense success rate of over $90\%$ against most attacks. In certain cases, MCGMark also experiences a failure in defense.
5.5 RQ4: Impact of Settings
To address RQ4, we employ the controlled variable method to investigate the impact of various display parameters on the watermark embedding success rate.
We explore the impact of three hyperparameters on the success rate of watermark embedding. We randomly select $50$ prompts from the $406$ tasks. Under the premise of unchanged settings elsewhere, we first investigate the impact of $\lceil\frac{|D|}{|V|}\rceil$ on the success rate of watermark embedding, where $\lceil\frac{|D|}{|V|}\rceil$ represents the ratio of vocabulary $\mathbb{A}$ or $\mathbb{B}$ in the total vocabulary during vocabulary partitioning. We fix the maximum output length of the LLM at $400$ and set $\lceil\frac{|D|}{|V|}\rceil$ to $[0.25,0.375,0.5,0.625,0.75]$, respectively. The setting of $\lceil\frac{|D|}{|V|}\rceil$ is based on the observed values of the literature [25]. Our results are shown in Fig. LABEL:fig:rq1. (a). The results show that the value of $\lceil\frac{|D|}{|V|}\rceil$ is not linearly related to the success rate of watermark embedding. In other words, increasing the vocabulary size does not necessarily make watermark embedding easier. In the current setting, $0.5$ appears to be the sweet spot value.
Furthermore, we explore the impact of the maximum number of output tokens of the LLM on watermark embedding. We kept the base setting of $\lceil\frac{|D|}{|V|}\rceil$. Simultaneously, we test the maximum number of output tokens by setting it to $[200,300,400,500,600]$, respectively. Our results are shown in Fig. LABEL:fig:rq1.(b). From the results, we can observe that as the maximum number of output tokens of the LLM increases, the success rate of watermark embedding also tends to increase. However, there is an evident diminishing marginal effect.
Finally, we explore the impact of hash value $H$ on the watermark Embedding Success Rate. In MCGMark, the hash value is continuously varied to ensure the randomness of vocabulary partitioning. Thus, for our purposes, we modify the hash strategy to a fixed approach, setting the hash value to $7,775$ and $666$, respectively. All other settings remained unchanged. Only nine watermarks were successfully embedded under the hash value of $7,775$, yielding an embedding success rate of $18\%$. Conversely, under the hash value of $666$, $19$ watermarks were successfully embedded, resulting in a success rate of $38\%$. We further test scenarios with $H$ values of $15,485,863$ and two, resulting in embedding success rates of $36\%$ and $44\%$, respectively. From this phenomenon, two conclusions can be drawn: (1) The choice of hash value significantly impacts the watermark embedding success rate, providing an opportunity for SSP to enhance performance by selecting appropriate hash values. (2) MCGMark’s pseudo-random hash strategy is effective and can substantially improve the watermark embedding success rate compared to fixed hash values. Utilizing MCGMark’s strategy, SSP no longer needs to expend additional time exploring hash values.
TABLE IV: The impact of fixed hash values on watermark embedding success rate.
Fixed Hash Key 7,775 666 15,485,865 2
Embedding Success Rate $18\%$ $38\%$ $36\%$ $44\%$
Answer to RQ 4: The maximum output token count of the LLM, the proportion of vocabulary partitioning $\lceil\frac{|D|}{|V|}\rceil$, and the hash value $H$ all influence the success rate of watermark embedding. This indirectly highlights the importance of mechanisms such as reproducible floating hashes that we have implemented.
6 Limitations and Threats to validity
6.1 Limitations
In Section 5.4, while the watermark demonstrates robust performance in most scenarios, it occasionally fails to defend against certain attacks. Our analysis of these failures reveals that incomplete token adaptation during watermark embedding is a primary issue. Currently, we rely on regular expressions for token recognition, but LLMs often have extensive vocabularies (3.2K tokens or more), making comprehensive adaptation challenging. In future work, we plan to develop a more robust token recognition approach or seek support from SSPs to better address this problem.
Another limitation is the impact of the watermark on the quality of LLM output. To address this, we design Algorithm 1 based on probabilistic outliers in the LLM’s vocabulary to ensure optimal token selection. However, the watermark inevitably impacts code quality. This occurs for two reasons: first, outliers themselves might have quality variations; second, the error correction bit of the watermark can influence the code output. Fortunately, the criteria for filtering outliers can be adjusted flexibly. A looser outlier filtering standard ensures more random model output, while a stricter standard guarantees higher output quality. Additionally, the introduction of the error correction bit reduces the watermark’s impact on model output by over $50\%$. Further compressing the length of the watermark, especially the error correction bit, can lead to even greater reductions in impact.
6.2 Threats to validity
6.2.1 Internal Validity
In section 3, the construction of the MCGTest relied on manual analysis by participants, which introduces a certain degree of subjectivity to the results. To mitigate this threat, we use a close card sorting method for each aspect requiring manual analysis, involving at least two participants to ensure consistency in the results. We make significant efforts to mitigate this potential threat.
Another internal validity concern in constructing MCGTest is our reliance on keyword matching for gathering code repositories and data, which may lead to incomplete data collection. However, the primary goal in creating MCGTest is to assemble a sufficient number of scenarios involving malicious code generation. This allows us to design, test, and inspire future developments in MCGMark. Our objective is not to capture every instance of malicious code but to collect representative cases that serve our research needs.
6.2.2 External Validity
We test MCGMark exclusively on DeepSeek-Coder-6.7b because it has demonstrated high performance with relatively small parameters in various benchmarks [5]. However, the watermark embedding, detection, and robustness algorithms proposed in this paper are not specific to any single LLM. They are designed to be applicable to any LLM based on the transformer architecture [78]. Adapting MCGMark to other models involves adjusting the code element matching in Algorithm 2 to fit the model’s vocabulary. This adjustment is straightforward and does not present significant technical challenges.
Similarly, this paper adapts watermark patterns specifically for the Python language, as it is currently the most commonly used language among attackers [79]. Adapting the watermark patterns to other languages is not a technical challenge; it merely requires adjusting and matching the code elements of the target language with the seven patterns of code elements. It is important to note that since our watermark is embedded online, we do not have access to the entire code during embedding. Therefore, we cannot extend MCGMark’s adaptability based on the Abstract Syntax Tree (AST).
7 Discussion
While evaluating MCGMark, we identify several scenarios that pose challenges to current watermarking techniques. This section discusses these scenarios and proposes potential solutions to address them.
7.1 Watermark Embedding Failure
Short Code Generation. Embedding multi-bit watermarks while ensuring robustness is particularly challenging in short code generation scenarios. Although this paper tests with 400 tokens (DeepSeek-Coder supports up to 2048 tokens, while models like Codellama generate sequences of up to 100,000 tokens), MCGMark struggles in extremely short code scenarios. Designing multi-bit watermarks for shortcode generation without compromising robustness remains a significant challenge. One potential solution involves compressing watermark encoding using higher-dimensional vocabulary partitions.
Poor Model Outputs. Watermark embedding often fails in scenarios where model outputs are subpar, such as generating code that doesn’t adhere to syntax rules or mixes natural language with code. Online watermark embedding heavily relies on the generation capabilities of LLMs. Transitioning to a more powerful model is the most direct solution. Additionally, embedding considerations should include rollback strategies based on tokens generated by LLMs to safeguard model outputs.
7.2 Watermark Detection Failure
Tokenization Discrepancies. Tokenizers are not always reliable. Minor discrepancies between tokenization during code generation and code detection may lead to errors in watermark detection. Notably, if tokens affected by these discrepancies do not contain watermarks, automatic correction may occur during subsequent detection processes. Conducting systematic empirical studies on tokenizer errors could provide insights into addressing this issue.
Watermark Strength. The strength of the watermark directly affects the success rate of watermark detection; lower strength results in detection failures, while excessively high strength impacts the output quality of the model. To mitigate this issue, this paper proposes Algorithm 1 and designs a watermark prompt. However, this approach still affects the output quality to some extent. Balancing watermark strength and detection success rate remains a significant challenge worth exploring.
8 Related work
8.1 Traditional Code Watermark
Code watermarking involves directly adding a special identifier to the source code or during code execution to declare code ownership [80]. Code watermarking techniques can be broadly categorized into static and dynamic approaches [19]. Static watermarking embeds watermarks directly into the source code. For instance, Kim et al. [81] uses adaptive semantic-preserving transformations to embed watermarks, and Sun et al. [82] does so by changing the order of functions. However, static watermarks are relatively more susceptible to detection and removal [83]. Consequently, dynamic code watermarking techniques have seen rapid development recently. For example, LLWM [84] is a watermarking technique that uses LLVM and Clang to embed watermarks by compiling the code. Xmark [85] embeds watermarks by obfuscating the control flow based on the Collatz conjecture. However, dynamic watermarking techniques are not applicable to the code generation process of LLMs, as LLMs do not execute the generated code.
Difference. Therefore, current traditional code watermarking techniques are not suitable for LLM code generation. The watermarking techniques mentioned above differ significantly from the watermark proposed in this paper, which operates during the LLM code generation process. We need to design more innovative watermarks for the LLM code generation task.
8.2 LLM Watermark
The security of LLMs has recently gained wide attention, leading to efforts to use watermarking techniques for their protection. Regarding watermarking techniques for model security, the aim is to prevent model theft through methods such as distillation. For example, GINSW [86] injects private signals into the probability vectors of the decoding step for each target token to prevent model theft. TOSYN [18] replaces code training data samples with synthesized code to protect the model against distillation attacks. PLMmark [87] ensures model ownership by watermarks in the LLM training.
Difference. The watermarking techniques proposed in above works can protect model copyright in various scenarios. However, they are not suitable for verifying text generated by LLMs. Therefore, there is a fundamental difference between these techniques and the problem addressed by the watermarking method proposed in this paper.
8.3 Watermarking for LLM Text Generation
Non-encodable Text Watermark. LLM watermarking techniques can be categorized into offline and online watermarking. In online watermarking, the watermarking process is synchronized with the LLM content generation process [88]. On the other hand, offline watermarking requires processing the generated text after the LLM has completed content generation [89, 90]. Offline watermarking is not directly related to the LLM itself. They rely more on rules and are easier for attackers to detect [20]. Kirchenbauer et al. [25] is a text watermarking technique that has received considerable attention. It involves partitioning the vocabulary of LLM and guiding the model to select words from a predefined vocabulary, thereby embedding watermarks into the generated text [88]. A follow-up work [88] investigates the reliability of this watermarking strategy. Subsequent works have extended this watermarking technique to privacy [91] and robustness [42].
Difference. However, this watermarking strategy can only obtain binary results. Recently, some research has focused on designing encodable watermarks for LLMs.
Encodable Text Watermark. Yoo et al. [43], also building on the work of Kirchenbauer et al. [25], achieved multi-bit watermark embedding by dividing the vocabulary into more sub-vocabularies. Wang et al. [92] analyzed how to embed more information into watermarks through vocabulary partitioning from a mathematical perspective. However, further subdividing the vocabulary may lead to a significant decline in the quality of the model’s output.
Difference. Therefore, the aforementioned works are only applicable to text-generation scenarios. Code, as a special type of text, has a relatively fixed structure and pattern. These watermarks are not designed for code generation scenarios and cannot address the issues proposed in this paper.
8.4 Watermarking for LLM Code Generation
As the issue of malicious developers using LLMs to generate malicious code becomes more recognized, some works have designed watermarking schemes specifically for LLM code generation. Li et al. [23] devised a set of code transformation rules to embed watermarks through post-processing. Yang et al. [24] designed an AST-based code transformation method to embed multi-bit watermarks, which is also a post-processing watermark. Post-processing watermarks are more susceptible to attackers discovering their rules and disrupting the watermark. Additionally, neither of these methods offers robust solutions tailored to the structure of the code. SWEET [21] is an online code watermark that extends the low-entropy scenarios of Kirchenbauer et al. [25] to improve code generation quality. However, SWEET can only achieve binary results and still does not address the issues proposed in this paper.
Difference. The approach proposed in this paper is fundamentally different from any of the above methods. It is an online watermarking scheme, making the watermark more challenging to detect and remove. Furthermore, our watermark is encodable and capable of embedding the creator’s identity information. It also enhances robustness against code structure to prevent malicious developers from easily breaking the watermark through simple modifications.
9 Conclusion
In this work, we propose MCGMark, a robust and encodable watermarking technique for LLMs to counteract the growing trend of malicious code generation. MCGMark leverages code structural features to embed user identity information while LLM token output. To minimize the impact on LLM code generation quality, we introduce a watermark quality enhancement scheme for MCGMark based on probabilistic outliers. Furthermore, we design a watermark skipping algorithm based on code structural features and syntax rules, excluding easily removable elements such as comments, thus enhancing MCGMark robustness. We also develop MCGTest, a dataset of $406$ malicious code generation tasks, derived from a review of existing malicious code. And we evaluate the effectiveness and robustness of MCGMark on MCGTest.
We make MCGTest and MCGMark open-source to inspire further countermeasures against malicious code generation, addressing the growing problem of malicious developers using LLMs to create harmful code. Additionally, in future work, we plan to explore watermark compression methods to reduce interference with code generation quality. We also intend to develop new token-matching techniques to accommodate the complex tokens in different LLMs, thereby enhancing the defense capabilities of the watermark.